# Energy solutions of the Cauchy–Dirichlet problem for fractional nonlinear diffusion equations

### v1

@@ 6.2. Quasi-convergence to asymptotic profiles
Theorem 6.2 (Łojasiewicz–Simon inequalities for $I$ (cf. [7, Theorem 5])).
@@
The case $\theta=1$ was first studied in [39]. The above-mentioned theorem is almost same as Theorem 5 of [7], whose proof is concerned with the fractional case $10<\theta<$ 1 but clearly available for the case $\theta=1$ as well. On the other hand, in (H1) and (H2) mentioned above, growth conditions are “local” in contrast with the original ones in [7, Theorem 5]; indeed, in (H1) above, the constants $C,M$ may depend on the choice of $\alpha,\beta$. However, this slight modification does not require any major change of the original proof in [7]. To be more precise, a simple modification is needed only for the proof of Proposition 6.2 of [7], where the analyticity around $\varphi$ of the operator $w\mapsto g(w)$ from $X^{\theta}_{p}:=\{w\in\mathcal{X}_{\theta}(\Omega)\cap L^{p}(\Omega)\colon(-\Delta)^{\theta}w\in L^{p}(\Omega)\}$ endowed with the norm $\|\cdot\|_{X^{\theta}_{p}}:=\|\cdot\|_{L^{p}(\Omega)}+\|\cdot\|_{\mathcal{X}_{\theta}(\Omega)}+\|(-\Delta)^{\theta}\cdot\|_{L^{p}(\Omega)}$ into $L^{p}(\Omega)$ is proved for $p\in(1,\infty)$ large enough. Indeed, as $d>2\theta$, the space $X^{\theta}_{p}$ is continuously embedded in $L^{\infty}(\Omega)$ for $p\in(1,\infty)$ large enough. [^2] In case $\theta=1$, it is nothing but the $L^{p}$ elliptic estimate and a Sobolev embedding. Indeed, $X^{\theta}_{p}$ coincides with $W^{2,p}(\Omega)\cap H^{1}_{0}(\Omega)$. Moreover, we have $W^{2,p}(\Omega)\hookrightarrow L^{\infty}(\Omega)$ for $p>d/2$ without any restriction of $d$ (cf. see [39]). On the other hand, such an $L^{p}$ elliptic estimate may fail for $10<\theta<$ 1. Hence any $w\in X^{\theta}_{p}$ lying on a small neighbourhood of $\varphi$ (in $X^{\theta}_{p}$) is uniformly bounded in $L^{\infty}(\Omega)$, say $\|w\|_{L^{\infty}(\Omega)}<r_{0}$ for some $r_{0}>0$. We then apply either (H1) or (H2) by taking $\alpha,\beta>r_{0}$ to prove the analyticity. ∎
@@ 6.4. Full-convergence to asymptotic profiles
First of all, assuming that $\Omega$ is a bounded $C^{1,1}$ domain of $\mathbb{R}^{d}$, one can prove $\phi\in L^{\infty}(\Omega)$ even for $10<\theta<$ 1 as in Appendix §D by performing a standard bootstrap argument (with the aid of (4.9)) and by exploiting a fractional elliptic regularity (see [60, Proposition 1.4]). The following lemma will be used for the fast diffusion case, i.e., $q>2$.
@@
Lemma 6.4 (Uniqueness of positive solutions for $21<q<$ 2 [19, Corollary 2.4][26, Theorem 1]).
@@
This lemma can be proved as in [19, Corollary 2.4]. Moreover, it also follows from a more general uniqueness result, which is a fractional variant of Brézis–Oswald’s uniqueness result [26] and which will be provided with a proof in Appendix §E.
@@
(see [2]). Since $J(v(\cdot))$ is nonincreasing on $[0,\infty)$ thanks to (i) of Lemma 6.1, so is $(J(v(\cdot))-J(\phi))^{\sigma}>0$. Hence recalling [8, Corollary 3.29], one can verify from (6.12) that
@@
In case $21<q<$ 2, suppose that $\phi\geq 0$ a.e. in $\Omega$. The full convergence (6.11) has already been proved in [19] by exploiting the Bénilan–Crandall estimate for (1.1)–(1.3) to guarantee the monotonicity of $v=v(x,s)$ in the rescaled time $s$. Here we can also prove (6.11) immediately from the quasi-convergence of $v$ (see Theorem 3.8) and the uniqueness of the positive solution to (3.26), (3.27) (see Lemma 6.4). This completes the proof. ∎
@@ Appendix B Density results in fractional Sobolev spaces
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ are the Fourier transform operator and its inverse, respectively, endowed with the norm $\|\cdot\|_{H^{\theta,p}(\mathbb{R}^{d})}$. In particular, we write $H^{\theta}(\mathbb{R}^{d})=H^{\theta,2}(\mathbb{R}^{d})$. It is shown in [56, Examples 1.8 and 2.12] that
@@
We refer the reader to [36, Theorems 2.4 and 5.4] for proofs of Theorems B.1 and B.2. In particular, the following theorem holds:
@@
We focus from now on to the case where $p=2$. As shown in [56, Corollary 4.37] (or [36, Proposition 3.4]), when $p=2$, the two definitions of fractional Sobolev spaces over $\mathbb{R}^{d}$ coincide, that is, $W^{\theta,2}(\mathbb{R}^{d})=H^{\theta}(\mathbb{R}^{d})$ with an equivalence of the norms. However, when $p\neq 2$, $H^{\theta,p}(\mathbb{R}^{d})$ is strictly included in $W^{\theta,p}(\mathbb{R}^{d})$, or the contrary, depending on the choice of $p$. This also justifies the following notation used hereafter:
@@
In what follows, we assume that $\Omega\subset\mathbb{R}^{d}$ is bounded and smooth (see (7.10) of [53, p. 34] for more detail) so that there exists a $C^{\infty}$-extension $\rho$ of the distance function $\mathrm{dist}(\cdot,\partial\Omega)$ to the boundary, i.e. $\rho\in C^{\infty}(\overline{\Omega})$, $\rho>0$ on $\Omega$, $\rho=0$ on $\partial\Omega$, and moreover,
@@
Moreover, $H^{\theta}_{0}(\Omega)$ is endowed with the induced norm of $H^{\theta}(\Omega)$, while $H^{1/2}_{00}(\Omega)$ is endowed with the norm $\|u\|_{H^{1/2}(\Omega)}+\|\rho^{-1/2}u\|_{L^{2}(\Omega)}$. The following theorems can be found in [53, Chap. 1, §9, §11]. We also refer the reader to [53, Chap. 1, §7] for the definition of the space $H^{\theta}(\partial\Omega)$, whenever $\Omega$ is a bounded open set of $\mathbb{R}^{d}$ with smooth boundary and $\theta\in(0,1)$.
Theorem B.4 ([53, Chap. 1, Theorem 9.4]).
@@
Theorem B.5 ([53, Chap. 1, Theorem 11.1]).
@@
Theorem B.6 ([53, Chap. 1, Theorem 11.4]).
@@
Theorem B.7 ([53, Chap. 1, Theorem 11.5]).
@@
Theorem B.8 ([53, Chap. 1, Theorem 11.7]).
@@
From Theorem B.9, we deduce that $\mathcal{X}_{\theta}(\Omega)=H^{\theta}_{0}(\Omega)=\overline{C_{c}^{\infty}(\Omega)}^{\|\cdot\|_{H^{\theta}(\Omega)}}$ as far as $\theta\neq 1/2$. On the other hand, by virtue of Theorems B.8 and B.9, we find that $\mathcal{X}_{1/2}(\Omega)=H^{1/2}_{00}(\Omega)=(L^{2}(\Omega),H^{1}_{0}(\Omega))_{1/2,2}$ with equivalence of norms. By general theory on real interpolation spaces (see [56, Proposition 1.17]), we conclude that
@@ Appendix C Regularization of initial data
and then, since $\phi\in L^{p_{n}}(\Omega)$, that is, $|\phi|^{q-2}\phi\in L^{p_{n}/(q-1)}(\Omega)$, due to [60, Proposition 1.4], since $\Omega$ is a bounded $C^{1,1}$ domain of $\mathbb{R}^{d}$, we can reach the Hölder regularity,
@@ Appendix E Uniqueness of positive solutions to fractional sublinear elliptic equations
Therefore (F1)–(F3) imply that the function $x\mapsto f(x,u(x))$ belongs to $L^{\infty}(\Omega)$. We deduce from the fractional elliptic regularity result [59, Proposition 1.4] that $u\in C^{\theta}(\mathbb{R}^{d})$. We can similarly show that $v\in C^{\theta}(\mathbb{R}^{d})$. Moreover, we see from (F1) and (F2) that there exists a constant $M>0$ such that

### v2

@@ 6.2. Quasi-convergence to asymptotic profiles
Theorem 6.2 (Łojasiewicz–Simon inequalities for $I$ (cf. [8, Theorem 5])).
@@
The case $\theta=1$ was first studied in [43]. The above-mentioned theorem is almost same as Theorem 5 of [8], whose proof is concerned with the fractional case $10<\theta<$ 1 but clearly available for the case $\theta=1$ as well. On the other hand, in (H1) and (H2) mentioned above, growth conditions are “local” in contrast with the original ones in [8, Theorem 5]; indeed, in (H1) above, the constants $C,M$ may depend on the choice of $\alpha,\beta$. However, this slight modification does not require any major change of the original proof in [8]. To be more precise, a simple modification is needed only for the proof of Proposition 6.2 of [8], where the analyticity around $\varphi$ of the operator $w\mapsto g(w)$ from $X^{\theta}_{p}:=\{w\in\mathcal{X}_{\theta}(\Omega)\cap L^{p}(\Omega)\colon(-\Delta)^{\theta}w\in L^{p}(\Omega)\}$ endowed with the norm $\|\cdot\|_{X^{\theta}_{p}}:=\|\cdot\|_{L^{p}(\Omega)}+\|\cdot\|_{\mathcal{X}_{\theta}(\Omega)}+\|(-\Delta)^{\theta}\cdot\|_{L^{p}(\Omega)}$ into $L^{p}(\Omega)$ is proved for $p\in(1,\infty)$ large enough. Indeed, as $d>2\theta$, the space $X^{\theta}_{p}$ is continuously embedded in $L^{\infty}(\Omega)$ for $p\in(1,\infty)$ large enough. [^2] In case $\theta=1$, it is nothing but the $L^{p}$ elliptic estimate and a Sobolev embedding. Indeed, $X^{\theta}_{p}$ coincides with $W^{2,p}(\Omega)\cap H^{1}_{0}(\Omega)$. Moreover, we have $W^{2,p}(\Omega)\hookrightarrow L^{\infty}(\Omega)$ for $p>d/2$ without any restriction of $d$ (cf. see [43]). On the other hand, such an $L^{p}$ elliptic estimate may fail for $10<\theta<$ 1. Hence any $w\in X^{\theta}_{p}$ lying on a small neighbourhood of $\varphi$ (in $X^{\theta}_{p}$) is uniformly bounded in $L^{\infty}(\Omega)$, say $\|w\|_{L^{\infty}(\Omega)}<r_{0}$ for some $r_{0}>0$. We then apply either (H1) or (H2) by taking $\alpha,\beta>r_{0}$ to prove the analyticity. ∎
@@ 6.4. Full-convergence to asymptotic profiles
First of all, assuming that $\Omega$ is a bounded $C^{1,1}$ domain of $\mathbb{R}^{d}$, one can prove $\phi\in L^{\infty}(\Omega)$ even for $10<\theta<$ 1 as in Appendix §D by performing a standard bootstrap argument (with the aid of (4.9)) and by exploiting a fractional elliptic regularity (see [65, Proposition 1.4]). The following lemma will be used for the fast diffusion case, i.e., $q>2$.
@@
Lemma 6.4 (Uniqueness of positive solutions for $21<q<$ 2 [29, Theorem 1][20, Corollary 2.4][44, Lemma 3.2]).
@@
This lemma can be proved as in [20, Corollary 2.4]. Moreover, it also follows from a more general uniqueness result, which is a fractional variant of Brézis–Oswald’s uniqueness result [29] and which will be provided with a proof in Appendix §E.
@@
(see [2]). Since $J(v(\cdot))$ is nonincreasing on $[0,\infty)$ thanks to (i) of Lemma 6.1, so is $(J(v(\cdot))-J(\phi))^{\sigma}>0$. Hence recalling [9, Corollary 3.29], one can verify from (6.12) that
@@
In case $21<q<$ 2, suppose that $\phi\geq 0$ a.e. in $\Omega$. The full convergence (6.11) has already been proved for nonnegative solutions in [20] by exploiting the Bénilan–Crandall estimate for (1.1)–(1.3) to guarantee the monotonicity of $v=v(x,s)\geq 0$ in the rescaled time $s$. Moreover, it is also verified in [44, Theorem 1.1] from a variational point of view for (possibly) sign-changing $v=v(x,s)$ emanating from initial data whose energies are less than the second least energy for (3.26), (3.27). Here we can also prove (6.11) immediately from the quasi-convergence of $v$ to a nonnegative asymptotic profile $\phi\geq 0$, which turns out to be positive in $\Omega$ due to the fractional strong maximum principle (see, e.g., [46, Theorem 2.1]), as well as the uniqueness of the positive solution to (3.26), (3.27) (see Lemma 6.4). This completes the proof. ∎
@@ Appendix B Density results in fractional Sobolev spaces
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ are the Fourier transform operator and its inverse, respectively, endowed with the norm $\|\cdot\|_{H^{\theta,p}(\mathbb{R}^{d})}$. In particular, we write $H^{\theta}(\mathbb{R}^{d})=H^{\theta,2}(\mathbb{R}^{d})$. It is shown in [61, Examples 1.8 and 2.12] that
@@
We refer the reader to [39, Theorems 2.4 and 5.4] for proofs of Theorems B.1 and B.2. In particular, the following theorem holds:
@@
We focus from now on to the case where $p=2$. As shown in [61, Corollary 4.37] (or [39, Proposition 3.4]), when $p=2$, the two definitions of fractional Sobolev spaces over $\mathbb{R}^{d}$ coincide, that is, $W^{\theta,2}(\mathbb{R}^{d})=H^{\theta}(\mathbb{R}^{d})$ with an equivalence of the norms. However, when $p\neq 2$, $H^{\theta,p}(\mathbb{R}^{d})$ is strictly included in $W^{\theta,p}(\mathbb{R}^{d})$, or the contrary, depending on the choice of $p$. This also justifies the following notation used hereafter:
@@
In what follows, we assume that $\Omega\subset\mathbb{R}^{d}$ is bounded and smooth (see (7.10) of [58, p. 34] for more detail) so that there exists a $C^{\infty}$-extension $\rho$ of the distance function $\mathrm{dist}(\cdot,\partial\Omega)$ to the boundary, i.e. $\rho\in C^{\infty}(\overline{\Omega})$, $\rho>0$ on $\Omega$, $\rho=0$ on $\partial\Omega$, and moreover,
@@
Moreover, $H^{\theta}_{0}(\Omega)$ is endowed with the induced norm of $H^{\theta}(\Omega)$, while $H^{1/2}_{00}(\Omega)$ is endowed with the norm $\|u\|_{H^{1/2}(\Omega)}+\|\rho^{-1/2}u\|_{L^{2}(\Omega)}$. The following theorems can be found in [58, Chap. 1, §9, §11]. We also refer the reader to [58, Chap. 1, §7] for the definition of the space $H^{\theta}(\partial\Omega)$, whenever $\Omega$ is a bounded open set of $\mathbb{R}^{d}$ with smooth boundary and $\theta\in(0,1)$.
Theorem B.4 ([58, Chap. 1, Theorem 9.4]).
@@
Theorem B.5 ([58, Chap. 1, Theorem 11.1]).
@@
Theorem B.6 ([58, Chap. 1, Theorem 11.4]).
@@
Theorem B.7 ([58, Chap. 1, Theorem 11.5]).
@@
Theorem B.8 ([58, Chap. 1, Theorem 11.7]).
@@
From Theorem B.9, we deduce that $\mathcal{X}_{\theta}(\Omega)=H^{\theta}_{0}(\Omega)=\overline{C_{c}^{\infty}(\Omega)}^{\|\cdot\|_{H^{\theta}(\Omega)}}$ as far as $\theta\neq 1/2$. On the other hand, by virtue of Theorems B.8 and B.9, we find that $\mathcal{X}_{1/2}(\Omega)=H^{1/2}_{00}(\Omega)=(L^{2}(\Omega),H^{1}_{0}(\Omega))_{1/2,2}$ with equivalence of norms. By general theory on real interpolation spaces (see [61, Proposition 1.17]), we conclude that
@@ Appendix C Regularization of initial data
and then, since $\phi\in L^{p_{n}}(\Omega)$, that is, $|\phi|^{q-2}\phi\in L^{p_{n}/(q-1)}(\Omega)$, due to [65, Proposition 1.4], since $\Omega$ is a bounded $C^{1,1}$ domain of $\mathbb{R}^{d}$, we can reach the Hölder regularity,
@@ Appendix E Uniqueness of positive solutions to fractional sublinear elliptic equations
Therefore (F1)–(F3) imply that the function $x\mapsto f(x,u(x))$ belongs to $L^{\infty}(\Omega)$. We deduce from the fractional elliptic regularity result [64, Proposition 1.4] that $u\in C^{\theta}(\mathbb{R}^{d})$. We can similarly show that $v\in C^{\theta}(\mathbb{R}^{d})$. Moreover, we see from (F1) and (F2) that there exists a constant $M>0$ such that